OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (zzr0ck3r):

-3|2x+6|=-12 3|2x+6|=12 |2x+6| = 4 so now we have two equations 2x+6 = 4 and 2x+6 = -4 so x = -1 and -5

OpenStudy (zzr0ck3r):

so what you want to do, is simplify until you get to |stuff| = something then stuff = something and stuff = -something

OpenStudy (zzr0ck3r):

@acama you understand?

OpenStudy (isaiah.feynman):

@acama zzr0ck3r is simply explaining this rule of absolute values.